# An LLM-Empowered Adaptive Evolutionary Algorithm For Multi-Component Deep Learning Systems

### v1

@@ Abstract
Multi-objective evolutionary algorithms (MOEAs) are widely used for searching optimal solutions in complex multi-component applications. Traditional MOEAs for multi-component deep learning (MCDL) systems face challenges in enhancing the search efficiency while maintaining the diversity. To combat these, this paper proposes $\mu$MOEA, the first LLM-empowered adaptive evolutionary search algorithm to detect safety violations in MCDL systems. Inspired by the context-understanding ability of Large Language Models (LLMs), $\mu$MOEA promotes the LLM to comprehend the optimization problem and generate an initial population tailed to evolutionary objectives. Subsequently, it employs adaptive selection and variation to iteratively produce offspring, balancing the evolutionary efficiency and diversity. During the evolutionary process, to navigate away from the local optima, $\mu$MOEA integrates the evolutionary experience back into the LLM. This utilization harnesses the LLM’s quantitative reasoning prowess to generate differential seeds, breaking away from current optimal solutions. We evaluate $\mu$MOEA in finding safety violations of MCDL systems, and compare its performance with state-of-the-art MOEA methods. Experimental results show that $\mu$MOEA can significantly improve the efficiency and diversity of the evolutionary search.
@@ Introduction
Table 1: Prompt patterns for LLM-based initial population
@@
(1) For each generation, $\mu$MOEA builds improved Pareto-optimal solutions considering multiple objectives to measure the potential of solutions to expose safety violations of the MCDL system. The criticality metric is used to evaluate how close the tested MCDL system’s behavior is to safety violations. The multi-objectives of $\mu$MOEA include criticality and diversity. Practically, criticality is formulated as: $f_{c}^{s_{i}}=\min_{t\in s_{i}}\{SV_{MCDL}^{t}\}$, where $SV$ is the distance to safety violations. The diversity metric is used to evaluate the coverage of the test cases generated by previous generations. Practically, diversity is formulated as: $f_{d}^{s_{i},s_{j}}=\frac{\sum_{n=1}^{|s_{i}|}\sum_{m=1}^{|s_{j}|}TD_{s_{i}^{n},s_{j}^{m}}}{|s_{i}|*|s_{j}|}$ The fitness function is represented as:
$S\leftarrow arg_{i\in G}\left\{minf_{c}(i),maxf_{d}(i)\right\}$
@@
$PM_{i}=\min\{k_{2}(f_{max}-f_{i})/(f_{max}-\overline{f}),k_{4}\},$
Table 2: Rules for feedback prompt generation
@@
We select the industrial full-stack ADS, Baidu Apollo (apo 2013) to evaluate the ability of $\mu$MOEA in finding safety violations of MCDL systems, due to the representativeness, practicality and advancedness. (1) Representativeness. Apollo ranks among the top 4 leading industrial ADS developers (Funicello-Paul April 1, 2024) (the other three ADSs, Waymo, Ford, and Cruise, are not released publicly). (2) Practicality. Apollo can be readily installed on vehicles for driving on public roads (Hersey April 1, 2024) (it has provided self-driving services for real vehicles (autowarefoundation Sepetem 1, 2023; Baidu April 1, 2024)). (3) Advancedness. Apollo is actively and rapidly updated (the releases of Apollo update on a weekly basis).
@@
To metric how distinct the different types of detected safety violations are, we calculate the average difference among them by Euclidean Distances. For two solutions $s_{i}$ and $s_{j}$, the calculation of Euclidean Distance, $ED_{s_{i},s_{j}}$ is given as follows:
$ED_{s_{i},s_{j}}=\frac{\sum_{n=1}^{|s_{i}|}\sum_{m=1}^{|s_{j}|}TD_{s_{i}^{n},s_{j}^{m}}}{|s_{i}|*|s_{j}|}$
The larger the distance, the more different the safety-violation types. $s_{i}^{n}$ represents the n-th chromosome in $s_{i}$. $|s_{i}|$ and $|s_{j}|$ represent the number of chromosomes in $s_{i}$ and $s_{j}$ respectively. $(x_{s_{i}^{n}.k},y_{s_{i}^{n}.k})$ represents the position of k-th gene of the n-th chromosome in $s_{i}$. $\alpha$ represents the minimal number of genes in the two chromosomes $s_{i}^{n}$ and $s_{j}^{m}$.
@@
Table 3: Comparison results of $\mu$MOEA, $\mu MOEA_{r}$ and $\mu MOEA_{n}$
$\mu$MOEA $\mu MOEA_{r}$ $\mu MOEA_{n}$
@@
(a) The found safety violation types over time
(b) The number of safety violations over time
Figure 2: The running results of $\mu$MOEA, $\mu MOEA_{r}$, $\mu MOEA_{n}$

### v2

@@ Abstract
Multi-objective evolutionary algorithms (MOEAs) are widely used for searching optimal solutions in complex multi-component applications. Traditional MOEAs for multi-component deep learning (MCDL) systems face challenges in enhancing the search efficiency while maintaining the diversity. To combat these, this paper proposes the first LLM-empowered adaptive evolutionary search algorithm to detect safety violations in MCDL systems. Inspired by the context-understanding ability of Large Language Models (LLMs), our approach promotes the LLM to comprehend the optimization problem and generate an initial population tailed to evolutionary objectives. Subsequently, it employs adaptive selection and variation to iteratively produce offspring, balancing the evolutionary efficiency and diversity. During the evolutionary process, to navigate away from the local optima, our approach integrates the evolutionary experience back into the LLM. This utilization harnesses the LLM’s quantitative reasoning prowess to generate differential seeds, breaking away from current optimal solutions. We evaluate our approach in finding safety violations of MCDL systems, and compare its performance with state-of-the-art MOEA methods. Experimental results show that our approach can significantly improve the efficiency and diversity of the evolutionary search.
@@ Introduction
Table 1: Prompt patterns for LLM-based initial population
@@
(1) For each generation, $\mu$MOEA builds improved Pareto-optimal solutions considering multiple objectives to measure the potential of solutions to expose safety violations of the MCDL system. The criticality metric is used to evaluate how close the tested MCDL system’s behavior is to safety violations. The multi-objectives of $\mu$MOEA include criticality and diversity. Practically, criticality is formulated as:
$f_{c}^{s_{i}}=\min_{t\in s_{i}}\{SV_{MCDL}^{t}\}$
where $SV$ is the distance to safety violations. The diversity metric is used to evaluate the coverage of the test cases generated by previous generations. Practically, diversity is formulated as:
$f_{d}^{s_{i}}=\frac{\sum_{j=1}^{\lvert PN\lvert}ED_{s_{i},s_{j}}}{\lvert PN\lvert},ED_{s_{i},s_{j}}=\frac{\sum_{n=1}^{|s_{i}|}\sum_{m=1}^{|s_{j}|}TD_{s_{i}^{n},s_{j}^{m}}}{|s_{i}|*|s_{j}|}$
The fitness function is represented as:
$S\leftarrow arg_{i\in G}\left\{minf_{c}^{s_{i}},maxf_{d}^{s_{i}}\right\}$
@@
$PM_{i}=\min\{k_{2}(f_{max}-f_{i})/(\overline{f}-f_{min}),k_{4}\},$
@@
We select the industrial full-stack ADS, Baidu Apollo (Baidu Apollo 2013) to evaluate the ability of $\mu$MOEA in finding safety violations of MCDL systems, due to the representativeness, practicality and advancedness. (1) Representativeness. Apollo ranks among the top 4 leading industrial ADS developers (Funicello-Paul April 1, 2024) (the other three ADSs, Waymo, Ford, and Cruise, are not released publicly). (2) Practicality. Apollo can be readily installed on vehicles for driving on public roads (Hersey April 1, 2024) (it has provided self-driving services for real vehicles (Autoware Foundation 2015; Baidu April 1, 2024)). (3) Advancedness. Apollo is actively and rapidly updated (the releases of Apollo update on a weekly basis).
@@
$\mu$MOEA $\mu$MOEAr $\mu$MOEAn
@@
Table 3: Results of $\mu$MOEA and ablation baselines
Figure 2: The number of found safety violations over time